# Equitable Allocation for Mixtures of Goods and Chores

### v2

@@ 3 EQ1 Allocations
When the input instance has subjective items, an EQ1 allocation may not exist, even for two agents and two items. The following example illustrates this.
Example 3.1.
Consider an instance with two agents (Alice and Bob) and two items such that Alice values the two items at $-1$ each while Bob values them at $1$ each. In any allocation, Alice’s valuation is strictly less than that of Bob even after the hypothetical removal of some item from any of the bundles. Therefore, no allocation in this instance is EQ1.
Moreover, we show in the following result that deciding if there is an EQ1 allocation is weakly NP-Hard. This is in contrast to the only goods or only chores setting where even for non-normalized instances, computing an EQ1 allocation admits simple and efficient algorithms (Gourvès et al., 2014; Freeman et al., 2020).
@@
For any allocation instance with mixed items, deciding the existence of an EQ1 allocation is NP-complete.
@@
We exhibit a reduction from $2$-partition, where given a multiset $U=\{b_{1},b_{2}\ldots b_{m}\}$ of integers with sum $2T$, the task is to decide if there is a partition into two subsets $S$ and $U\setminus S$ such that sum of the numbers in both the partitions equals $T$. We construct the reduced allocation instance as follows. We create $4$ agents, $m$ set-items $\{o_{1},\ldots,o_{m}\}$, and $4$ dummy items $\{d_{1},\ldots d_{4}\}$. The first two agents value the set-items at $\{b_{1},\ldots b_{m}\}$ and all the dummy items at $-3T$. The last two agents value the set items at 00 and the dummy items at $T$. This completes the construction (see Table 2).
$o_{1}$ $o_{2}$ldots $o_{m}$ $d_{1}$ $d_{2}$ $d_{3}$ $d_{4}$
$a_{1}$ $b_{1}$ $b_{2}$ldots $b_{m}$ $-3T$ $-3T$ $-3T$ $-3T$
$a_{2}$ $b_{1}$ $b_{2}$ldots $b_{m}$ $-3T$ $-3T$ $-3T$ $-3T$
$a_{3}$ 00 00 $\ldots$ 00 $T$ $T$ $T$ $T$
$a_{4}$ 00 00 $\ldots$ 00 $T$ $T$ $T$ $T$
@@
The equivalence proof is relegated to the appendix. ∎
@@ 3.1 Objective Valuations
Despite the negative existence and computational results, we identify tractable instances for which an EQ1 allocation always exist and can be computed efficiently. In particular, we develop an algorithm that computes an EQ1 allocation when all items are either objective goods or objective chores. This finding gives rise to an intriguing relationship between partial and complete EQ1 allocations, as we state in Lemma 3.5.
@@
For any mixed instance with objective valuations, an EQ1 allocation always exists and can be computed in polynomial time.
@@
The algorithm iteratively picks a poor agent $p$—breaking ties arbitrarily—and allocates the most valuable good among the remaining items in $O^{+}$ according to $p^{\prime}$ valuation. Once $O^{+}$ is exhausted, it picks a rich agent $r$, who then receives its most disliked chore from $O^{-}$. The correctness argument is deferred to the appendix. ∎
@@
Consider the following instance. The output of the algorithm in Proposition 3.3 is highlighted, which fails EQX. Indeed, if Alice chooses to ignore the item $o_{2}$ from Bob’s bundle and hypothetically reduces its utility from $1$ to $-1$, it still remains a poor agent. But this allocation is EQ1, since Alice can ignore the chore $o_{7}$ from its bundle, thereby deriving equal utility as Bob, that is, $1$.
@@ 4 Normalized Valuations
In the quest for tractability, we now turn our attention to subjective but normalized and type-normalized valuations.
4.1 Symmetric Normalized Valuations
@@
For any mixed instance with $\{-1,1\}$ normalized valuations, an EQ1 allocation always exists and can be computed in polynomial time.
Note that since the instance is normalized, the reduced instance restricted to items in only $O^{\pm}$ is also normalized. That is, every agent assigns a value of $1$ to exactly $k_{1}$ items from $O^{\pm}$ and a value of $-1$ to exactly $k_{2}$ items from $O^{\pm}$, where $k_{1}$ and $k_{2}$ are constants. We show that there is a partial EQ1 allocation that allocates all the items in $O^{\pm}$ and then extend it to a complete allocation using Lemma 3.5.
@@
Output: An EQ1 allocation $A$
1:  $A\leftarrow$ An empty allocation
2:  $P\leftarrow$ Set of all poor agents under $A$
3:  $R\leftarrow$ Set of all rich agents under $A$
@@
5:     while Any $p$ values an unallocated item $o$ in $O^{\pm}$ at $1$ do
@@
8:     while Any $r$ values an unallocated item $o$ in $O^{\pm}$ at $-1$ do
@@
11:     Execute any of the feasible transfers described in the points (1-4)
@@
13:     if none of the transfers from (1-4) is feasible then
@@
26:  return   $A$
To that end, we first describe Algorithm 1. We iteratively pick a poor agent and allocate it an item it values at $1$ (a good) from $O^{\pm}$. If there are no such items in $O^{\pm}$, then we iteratively pick one of the rich agents and allocate it an item it values at $-1$ (a chore) from $O^{\pm}$. At this point, it is easy to see that the partial allocation is EQ1. We denote the set of agents with minimum utility as poor agents $(P)$ and those with maximum utility as rich agents ($R$). Now suppose that the remaining items in $O^{\pm}$ are all chores for the poor agents and goods for the rich agents, and if they are allocated to any of them, they increase the amount of inequity and hence may violate EQ1. Therefore, in order to move towards an EQ1 allocation that allocates all the items in $O^{\pm}$, we now aim to convert a poor (rich) agent to rich (poor) by re-allocating one of the allocated items, so that the converted agent can now be a potential owner of one of the remaining unallocated items while maintaining EQ1. We do so by executing one of the following transfers at a time, each of which allows us to allocate at least one unallocated item while respecting EQ1.
@@
    Rich-Rich Transfer (R-R). If $\exists\leavevmode\nobreak\ r,r^{\prime}\in R\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ o\in A_{r}:v_{r}(o)=1\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ v_{r^{\prime}}(o)=-1$, transfer $o$ from $A_{r}$ to $A_{r^{\prime}}$.
@@
    Rich-Poor Transfer (R-P). If $\exists\leavevmode\nobreak\ r\in R,p\in P\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ o\in A_{r}:v_{r}(o)=1=v_{p}(o)$, transfer $o$ from $A_{r}$ to $A_{p}$.
@@
    Poor-Rich Transfer (P-R). If $\exists\leavevmode\nobreak\ r\in R,p\in P\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ o\in A_{p}:v_{p}(o)=-1=v_{r}(o)$, transfer $o$ from $A_{p}$ to $A_{r}$.
@@
    Poor-Poor Transfer (P-P). If $\exists\leavevmode\nobreak\ p,p^{\prime}\in P\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ o\in A_{p}:v_{p}(o)=-1\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ v_{p^{\prime}}(o)=1$, transfer $o$ from $A_{p}$ to $A_{p^{\prime}}$.
@@
We show that Algorithm 1 returns an EQ1 allocation. Suppose there are no feasible transfers. If $|P|>|R|$, then we take a good ($1$-valued item) from $A_{r}$ and allocate it to $p$ who necessarily values it at $-1$ (otherwise, we would have implemented R-P transfer). We do this for $|R|$ disjoint pairs of a rich and a poor agent, thereby decreasing the utility of all these pairs by $-1$. But since $|P|>|R|$, we have a poor agent $p^{\prime}$ whose value remains intact, which in turn, makes him one of the rich agents after the above transfers. Now, the remaining unallocated item, which was a chore for $p^{\prime}$ can be allocated to it without violating EQ1. When $|P|\leq|R|$, then we argue that no feasible transfers imply a contradiction by using the assumption of normalization. We defer this case to the appendix. ∎
It is easy to see that the output allocation $A$ of Algorithm 1 is not only EQ1 but also EQX. Since all the goods in a rich agent’s bundle are valued at $1$ and all the chores in a poor agent’s bundle are valued at $-1$, removal of any single item gets rid of the inequity. This gives us the following result.
Corollary 4.2.
For any mixed instance with $\{-1,1\}$ normalized valuations, an EQX allocation always exists and can be computed in polynomial time.
If the instances are allowed to have 00-valued items, then the approach of assigning $O^{\pm}$ first and then completing the allocation fails even for type-normalized valuations, as illustrated by the following example. Note that in this case, $O\pm$ may not be normalized even if the entire instance is normalized.
@@
Alice $1$ $1$ $1$ 00 00 00 $-1$ $-1$ $-1$
Bob $-1$ $-1$ $-1$ $1$ $1$ $1$ 00 00 00
Therefore, when we have $\{-1,0,1\}$ valuations, we need a slightly different approach of allocating items from the entire bundle rather than exhausting a subset of items first. Using this approach, we present an efficient algorithm for this case.
@@
For any mixed instance with $\{-1,0,1\}$ normalized valuations, an EQ1 allocation always exists and can be computed in polynomial time.
The proof idea is similar to that of Theorem 4.1. We continue with the greedy algorithm until every poor agent values the remaining items at $-1$ and every rich agent values them at $1$. We then execute the transfers described below. The claim is that either we arrive at an EQ1 allocation or we arrive at a contradiction by leveraging the assumption of normalization. We defer the complete proof to the appendix, but present the set of feasible transfers below. After one of these transfers is implemented, at least one of the involved agents becomes a potential owner of an unallocated item without violating EQ1.
@@
Note that the above result does not show the existence of EQX allocation, but does show the existence of a weaker notion of ${}^{+}_{-}$ allocation Zhou et al. (2024), that entails the removal of items with non-zero marginals, that is, either a strictly positive valued item or a strictly negative valued item.
4.2 Challenges Beyond {-1, 0, 1} Valuations
@@
In addition, a leximin++ allocation (defined below), which is known to satisfy not only EQ1 but stronger EQX property for certain restricted instances, fails to achieve EQ1 for mixtures, even with normalized valuations. A leximin allocation is one that maximizes the minimum utility of an agent, subject to that maximizes the second-minimum utility, and so forth. Although there can be many leximin optimal allocations, they all induce a unique utility vector. Leximin++ allocation is an allocation that maximizes the minimum utility, and then maximizes the size of the bundle of the agent with the minimum utility, and so on. It is known that a leximin++ allocation is also EQX, for objective valuations with identical chores (Barman et al., 2024). When we go beyond objective valuations, Example 4.5 shows that leximin++ may not be EQ1, even for a normalized instance with two agents and two items. Additionally, Cousins et al. (2024) showed that for $\{-1,0,w\}$ (where $w$ is a positive integer) valuations, leximin allocations are EF1. However, for $\{-1,1\}$ valuations, even leximin++ may violate EQ1 as illustrated by Example A.2 in the appendix.
@@
Leximin++ is not EQ1 even for a normalized instance with two agents $(a_{1},a_{2})$ and two items $(o_{1},o_{2})$. $a_{1}$ values the items at $(10,-15)$ while $a_{2}^{\prime}s$ valuation is $(-2,-3)$. There are at most $4$ possible allocations in this instance with utility vectors $(0,-5),(-5,0),(10,-3),(-15,-2)$ for $(a_{1},a_{2})$ respectively. The only leximin++ allocation is the one with utility vector $(10,-3)$ but it is not EQ1. Even if we restrict the valuations to $(1,-1)$, Example $7$ (in the appendix) suggests that leximin++ may not be EQ1.

### v3

@@ 3 EQ1 Allocations
When the input instance has subjective items, an EQ1 allocation may not exist, even for two agents and two items, as illustrated by Example 1.1. Moreover, we show in the following result that deciding if there is an EQ1 allocation is weakly NP-Hard. This is in contrast to the only goods or only chores setting where even for non-normalized instances, computing an EQ1 allocation admits simple and efficient algorithms (Gourvès et al., 2014; Freeman et al., 2020).
@@
For any mixed fair division instance, deciding the existence of an EQ1 allocation is (weakly) NP-complete.
@@
We exhibit a reduction from $2$-partition, where given a multiset $U=\{b_{1},b_{2}\ldots b_{m}\}$ of integers with sum $2T$, the task is to decide if there is a partition into two subsets $S$ and $U\setminus S$ such that sum of the numbers in both the partitions equals $T$. We construct the reduced allocation instance as follows. We create $4$ agents, $m$ set-items $\{o_{1},\ldots,o_{m}\}$, and $4$ dummy items $\{d_{1},\ldots d_{4}\}$. The first two agents value the set-items at $\{b_{1},\ldots b_{m}\}$ and all the dummy items at 3T-3T. The last two agents value the set items at 0 and the dummy items at $T$. This completes the construction (see Table 2).
$o_{1}$ $o_{2}$ $\ldots$ $o_{m}$ $d_{1}$ $d_{2}$ $d_{3}$ $d_{4}$
$a_{1}$ $b_{1}$ $b_{2}$ $\ldots$ $b_{m}$ 3T-3T 3T-3T 3T-3T 3T-3T
$a_{2}$ $b_{1}$ $b_{2}$ $\ldots$ $b_{m}$ 3T-3T 3T-3T 3T-3T 3T-3T
$a_{3}$ 0 0 $\ldots$ 0 $T$ $T$ $T$ $T$
$a_{4}$ 0 0 $\ldots$ 0 $T$ $T$ $T$ $T$
@@
We now argue the equivalence.
Forward Direction.
Reverse Direction.
Suppose there is an EQ1 allocation, say $A$, under the reduced instance. Then, note that $A$ assigns at most one dummy item to the agents $a_{1}$ and $a_{2}$, each. If not, say WLOG, $a_{1}$ receives $\{d_{1},d_{2}\}$ under $A$, then the maximum utility $a_{1}$ can derive is 4T-4T (where it gets all the set items as well). But, $a_{1}$ violates EQ1 with respect to $a_{3}$ and $a_{4}$, whose minimum utility is 0 each, and $a_{1}$ derives negative utility even if it removes the item $d_{1}$ from its bundle. Therefore, $a_{1}$ and $a_{2}$ can get at most one dummy item each. We now consider the following cases:
  • $a_{1}$ receives $d_{1}$ and $a_{2}$ does not receive any dummy item. Then, WLOG, $a_{3}$ gets one dummy item and $a_{4}$ gets two dummy items. To be consistent with EQ1 against $a_{4}$, agent $a_{1}$ must get all the set items, thereby deriving a utility of 3T+2T-3T+2T. This leaves $a_{2}$ empty-handed and hence, it violates EQ1 with respect to $a_{4}$. Therefore, since $A$ is EQ1, this case does not arise.
This concludes the argument. ∎
@@ 3.1 Objective Valuations
Given the negative existence and computational results, we identify tractable instances for which an EQ1 allocation always exist and can be computed efficiently. In particular, we develop an algorithm that computes an EQ1 allocation when all items are either objective goods or objective chores. This finding gives rise to an intriguing relationship between partial and complete EQ1 allocations, as we state in Lemma 3.4.
@@
For any mixed fair division instance with objective valuations, an EQ1 allocation always exists and can be computed in polynomial time.
@@
The algorithm iteratively picks a poor agent pp—breaking ties arbitrarily—and allocates the most valuable good among the remaining items in $O^{+}$ according to $p$’s valuation. Once $O^{+}$ is exhausted, it picks a rich agent $r$, who then receives its most disliked chore from $O^{-}$. The correctness argument is deferred to the appendix. ∎
@@
Consider the following instance. The output of the algorithm in Proposition 3.2 is highlighted, which fails EQX. Indeed, if Alice chooses to ignore the item $o_{2}$ from Bob’s bundle and hypothetically reduces its utility from $1$ to 1-1, it still remains a poor agent. But this allocation is EQ1, since Alice can ignore the chore $o_{7}$ from its bundle, thereby deriving equal utility as Bob, that is, $1$.
@@ 4 Normalized Valuations
In the further quest for tractability, we now turn our attention to subjective but normalized and type-normalized valuations.
4.1 Symmetric Valuations
@@
For any mixed fair division instance with symmetric bi-valued normalized valuations ($\{-1,1\}$), an EQX (hence, EQ1) allocation always exists and can be computed in polynomial time.
It is easy to see that for $\{-1,1\}$ valuations, EQ1 and EQX allocations coincide. Since all the goods in a rich agent’s bundle are valued at $1$ and all the chores in a poor agent’s bundle are valued at 1-1, removal of any single item gets rid of the inequity. Therefore, it is sufficient to establish Theorem 4.1 for EQ1 allocations.
Note that since the instance is normalized, the reduced instance restricted to items in only $O^{\pm}$ is also normalized. That is, every agent assigns a value of $1$ to exactly $k_{1}$ items from $O^{\pm}$ and a value of 1-1 to exactly $k_{2}$ items from $O^{\pm}$, where $k_{1}$ and $k_{2}$ are constants. We show that there is a partial EQ1 allocation that allocates all the items in $O^{\pm}$ and then extend it to a complete allocation using Lemma 3.4.
@@
Output: An EQ1 allocation AA
[…]
@@
5:  while Any $p$ values an unallocated item $o$ in $O^{\pm}$ at 11 do
@@
8:  while Any $r$ values an unallocated item $o$ in $O^{\pm}$ at 1-1 do
@@
11:   Execute any of the feasible transfers described in the points (1-4) in Section 4.1
@@
13:  if none of the transfers in Step 11 is feasible then
@@
26:return  AA
To that end, we first describe Algorithm 1. We iteratively pick a poor agent and allocate it an item it values at $1$ (a good) from $O^{\pm}$. If there are no such items in $O^{\pm}$, then we iteratively pick one of the rich agents and allocate it an item it values at 1-1 (a chore) from $O^{\pm}$. At this point, it is easy to see that the partial allocation is EQ1. We denote the set of agents with minimum utility as poor agents $(P)$ and those with maximum utility as rich agents ($R$). Now suppose that the remaining items in $O^{\pm}$ are all chores for the poor agents and goods for the rich agents, and if they are allocated to any of them, they increase the amount of inequity and hence may violate EQ1. Therefore, in order to move towards an EQ1 allocation that allocates all the items in $O^{\pm}$, we now aim to convert a poor (rich) agent to rich (poor) by re-allocating one of the allocated items, so that the converted agent can now be a potential owner of one of the remaining unallocated items while maintaining EQ1. We do so by executing one of the following transfers at a time, each of which allows us to allocate at least one unallocated item while respecting EQ1.
@@
    Rich-Rich Transfer (R-R). If there exist agents $r,r^{\prime}\in R\leavevmode\nobreak\ \text{and an item}\leavevmode\nobreak\ o\in A_{r}\leavevmode\nobreak\ \text{such that}\leavevmode\nobreak\ v_{r}(o)=1\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ v_{r^{\prime}}(o)=-1$, transfer $o$ from $A_{r}$ to $A_{r^{\prime}}$.
@@
    Rich-Poor Transfer (R-P). If there exist agents $r\in R,p\in P\leavevmode\nobreak\ \text{and an item}\leavevmode\nobreak\ o\in A_{r}\leavevmode\nobreak\ \text{such that}\leavevmode\nobreak\ v_{r}(o)=1=v_{p}(o)$, transfer $o$ from $A_{r}$ to $A_{p}$.
@@
    Poor-Rich Transfer (P-R). If there exist agents $r\in R,p\in P\leavevmode\nobreak\ \text{and an item}\leavevmode\nobreak\ o\in A_{p}\leavevmode\nobreak\ \text{such that}\leavevmode\nobreak\ v_{p}(o)=-1=v_{r}(o)$, transfer $o$ from $A_{p}$ to $A_{r}$.
@@
    Poor-Poor Transfer (P-P). If there exist agents $p,p^{\prime}\in P\leavevmode\nobreak\ \text{and an item}\leavevmode\nobreak\ o\in A_{p}\leavevmode\nobreak\ \text{such that}\leavevmode\nobreak\ v_{p}(o)=-1\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ v_{p^{\prime}}(o)=1$, transfer $o$ from $A_{p}$ to $A_{p^{\prime}}$.
@@
We show that Algorithm 1 returns an EQ1 allocation. Suppose there are no feasible transfers. If $|P|>|R|$, then we take a good ($1$-valued item) from $A_{r}$ and allocate it to $p$ who necessarily values it at 1-1 (otherwise, we would have implemented R-P transfer). We do this for $|R|$ disjoint pairs of a rich and a poor agent, thereby decreasing the utility of all these pairs by 1-1. But since $|P|>|R|$, we have a poor agent $p^{\prime}$ whose value remains intact, which in turn, makes him one of the rich agents after the above transfers. Now, the remaining unallocated item, which was a chore for $p^{\prime}$ can be allocated to it without violating EQ1. When $|P|\leq|R|$, then we argue that no feasible transfers imply a contradiction by using the assumption of normalization. We defer this case to the appendix. ∎
If the instances are allowed to have 0-valued items, then the approach of assigning $O^{\pm}$ first and then completing the allocation fails even for type-normalized valuations, as illustrated by the following example. Note that in this case, $O\pm$ may not be normalized even if the entire instance is normalized.
@@
Alice 11 11 11 0 0 0 1-1 1-1 1-1
Bob 1-1 1-1 1-1 11 11 11 0 0 0
Therefore, when we have $\{-1,0,1\}$ valuations, we need a slightly different approach of allocating items from the entire bundle rather than exhausting a subset of items first. Using this approach, we present an efficient algorithm for finding EQ1 allocations. It is relevant to note here that in the presence of 0 valued items, EQ1 allocations do not coincide with EQX allocations, since removal of a good valued at 0 from the richer bundle is not helpful towards getting rid of the inequity.
@@
For any mixed fair division instance with symmetric tri-valued normalized valuations ($\{-1,0,1\}$), an EQ1 allocation always exists and can be computed in polynomial time.
The proof idea is similar to that of Theorem 4.1. We continue with the greedy algorithm until every poor agent values the remaining items at 1-1 and every rich agent values them at $1$. We then execute the transfers described below. The claim is that either we arrive at an EQ1 allocation or we arrive at a contradiction by leveraging the assumption of normalization. We defer the complete proof to the appendix, but present the set of feasible transfers below. After one of these transfers is implemented, at least one of the involved agents becomes a potential owner of an unallocated item without violating EQ1.
@@
    Rich-Rich Transfer (R-R). If there exist agents $r,r^{\prime}\in R\leavevmode\nobreak\ \text{and an item}\leavevmode\nobreak\ o\in A_{r}\leavevmode\nobreak\ \text{such that}\leavevmode\nobreak\ (v_{r}(o),v_{r^{\prime}}(o))\in\{(1,0),(0,-1),(1,-1)\}$, transfer $o$ from $A_{r}$ to $A_{r^{\prime}}$.
@@
    Rich-Poor Transfer (R-P). If there exist agents $r\in R,p\in P\leavevmode\nobreak\ \text{and an item}\leavevmode\nobreak\ o\in A_{r}\leavevmode\nobreak\ \text{such that}\leavevmode\nobreak\ (v_{r}(o),v_{p}(o))\in\{(1,1),(1,0),(0,1)\}$, transfer $o$ from $A_{r}$ to $A_{p}$.
@@
    Poor-Rich Transfer (P-R). If there exist agents $r\in R,p\in P\leavevmode\nobreak\ \text{and an item}\leavevmode\nobreak\ o\in A_{p}\leavevmode\nobreak\ \text{such that}\leavevmode\nobreak\ (v_{p}(o),v_{r}(o))\in\{(-1,0),(-1,-1),(0,-1)\}$, transfer $o$ from $A_{p}$ to $A_{r}$.
@@
    Poor-Poor Transfer (P-P). If there exist agents $p,p^{\prime}\in P\leavevmode\nobreak\ \text{and an item}\leavevmode\nobreak\ o\in A_{p}\leavevmode\nobreak\ \text{such that}\leavevmode\nobreak\ (v_{p}(o),v_{p^{\prime}}(o))\in\{(-1,1),(0,1),(-1,0)\}$, transfer $o$ from $A_{p}$ to $A_{p^{\prime}}$.
[…]
4.2 Challenges Beyond Symmetric Tri-valued Valuations
@@
In addition, a leximin++ allocation (defined below), which is known to satisfy not only EQ1 but stronger EQX property for certain restricted instances, fails to achieve EQ1 for mixtures, even with normalized valuations. A leximin allocation is one that maximizes the minimum utility of an agent, subject to that maximizes the second-minimum utility, and so forth. Although there can be many leximin optimal allocations, they all induce a unique utility vector. Leximin++ allocation is an allocation that maximizes the minimum utility, and then maximizes the size of the bundle of the agent with the minimum utility, and so on. It is known that a leximin++ allocation is also EQX, for objective valuations with identical chores (Barman et al., 2024). When we go beyond objective valuations, Example 4.4 shows that leximin++ may not be EQ1, even for a normalized instance with two agents and two items. Additionally, Cousins et al. (2024) showed that for $\{-1,0,w\}$ (where $w$ is a positive integer) valuations, leximin allocations are EF1.
@@
Leximin++ is not EQ1 even for a normalized instance with two agents $(a_{1},a_{2})$ and two items $(o_{1},o_{2})$. $a_{1}$ values the items at (10,15)(10,-15) while $a_{2}^{\prime}s$ valuation is (2,3)(-2,-3). There are at most $4$ possible allocations in this instance with utility vectors (0,5),(5,0),(10,3),(15,2)(0,-5),(-5,0),(10,-3),(-15,-2) for $(a_{1},a_{2})$ respectively. The only leximin++ allocation is the one with utility vector (10,3)(10,-3), but it is not EQ1.
Even if we restrict the valuations to (1,1)(1,-1), Example 4.5 suggests that leximin++ may not be EQ1.
Example 4.5.
Leximin++ is not EQ1 even with $\{1,-1\}$-normalized valuations. The instance in Table 4 illustrates this. The highlighted allocation is leximin++ with the utility vector as $\{3,1,1,1,1\}$. This is not EQ1 since even if the last $5$ agents choose to hypothetically ignore a good from $a_{1}^{\prime}s$ bundle, they still fall short of the equitable utility.
$o_{1}$ $o_{2}$ $o_{3}$ $o_{4}$ $o_{5}$ $o_{6}$ $o_{7}$ $o_{8}$ $o_{9}$ $o_{10}$ $o_{12}$ $o_{13}$
$a_{1}$ 1-1 11 11 11 11 11 -1 -1 1-1 1-1 1-1 1-1
$a_{2}$ 11 1-1 1-1 1-1 1-1 1-1 1-1 1-1 11 11 11 11
$a_{3}$ 11 1-1 1-1 1-1 1-1 1-1 1-1 1-1 11 11 11 11
$a_{4}$ 11 1-1 1-1 1-1 1-1 1-1 1-1 1-1 11 11 11 11
$a_{5}$ 11 1-1 1-1 1-1 1-1 1-1 1-1 1-1 11 11 11 11
$a_{6}$ 11 1-1 1-1 1-1 1-1 1-1 1-1 1-1 11 11 11 11